# Egor Cherepanov1,2Nikita Kachaev1Alexey K. Kovalev1,2Aleksandr I. Panov1,2 1AIRI, Moscow, Russia  2MIPT, Dolgoprudny, Russia

### v1

@@ F.10 MemoryGym Mystery Path
Mystery Path (Pleines et al., 2023) presents an invisible pathway that the agent must traverse without deviating. If the agent steps off the path, it is returned to the starting position, forcing it to remember the correct trajectory. The path is procedurally generated, meaning each episode introduces a new configuration. Success in this environment requires the agent to accurately recall previous steps and missteps to avoid repeating errors. The agent is rewarded +0.1 for progressing onto a previously unvisited tile
F.11 POPGym environments
@@
The following environments are included from the POPGym benchmark (Morad et al., 2023), which is designed to evaluate RL agents in partially observable settings. POPGym provides a diverse collection of lightweight vectorized environments with varying difficulty levels.
F.11.1 POPGym Autoencode
@@ F.11.2 POPGym Concentration
F.11.3 POPGym Repeat First
@@
F.11.4 POPGym Repeat Previous
@@
F.11.5 POPGym Stateless Cartpole
This is a modified version of the traditional Cartpole environment (Barto et al., 1983) where angular and linear position information is removed from observations. Instead, the agent only receives velocity-based data and must infer positional states by integrating this information over time to successfully balance the pole.
F.11.6 POPGym Stateless Pendulum
In this variation of the swing-up pendulum environment (Doya, 1995), angular position data is omitted from the agent’s observations. The agent must infer the pendulum’s position by processing velocity information and use this estimate to determine appropriate control actions.
F.11.7 POPGym Noisy Stateless Cartpole
@@
F.11.8 POPGym Noisy Stateless Pendulum
@@
F.11.9 POPGym Multiarmed Bandit
The Multiarmed Bandit environment is an episodic formulation of the multiarmed bandit problem (Slivkins, 2024), where a set of bandits is randomly initialized at the start of each episode. Unlike conventional multiarmed bandit tasks, where reward probabilities remain fixed across episodes, this structure resets them every time. The agent must dynamically adjust its exploration and exploitation strategies to maximize long-term rewards.
F.11.10 POPGym Higher Lower
@@
F.11.11 POPGym Count Recall
@@ F.11.12 POPGym Battleship
F.11.13 POPGym Mine Sweeper
@@
F.11.14 POPGym Labyrinth Explore
@@
F.11.15 POPGym Labyrinth Escape

### v2

@@ I.10 MemoryGym Mystery Path
Mystery Path [75] presents an invisible pathway that the agent must traverse without deviating. If the agent steps off the path, it is returned to the starting position, forcing it to remember the correct trajectory. The path is procedurally generated, meaning each episode introduces a new configuration. Success in this environment requires the agent to accurately recall previous steps and missteps to avoid repeating errors. The agent is rewarded +0.1 for progressing onto a previously unvisited tile
I.11 POPGym environments
@@
The following environments are included from the POPGym benchmark [65], which is designed to evaluate RL agents in partially observable settings. POPGym provides a diverse collection of lightweight vectorized environments with varying difficulty levels.
I.11.1 POPGym Autoencode
@@ I.11.2 POPGym Concentration
I.11.3 POPGym Repeat First
@@
I.11.4 POPGym Repeat Previous
@@
I.11.5 POPGym Stateless Cartpole
This is a modified version of the traditional Cartpole environment [6] where angular and linear position information is removed from observations. Instead, the agent only receives velocity-based data and must infer positional states by integrating this information over time to successfully balance the pole.
I.11.6 POPGym Stateless Pendulum
In this variation of the swing-up pendulum environment [18], angular position data is omitted from the agent’s observations. The agent must infer the pendulum’s position by processing velocity information and use this estimate to determine appropriate control actions.
I.11.7 POPGym Noisy Stateless Cartpole
@@
I.11.8 POPGym Noisy Stateless Pendulum
@@
I.11.9 POPGym Multiarmed Bandit
The Multiarmed Bandit environment is an episodic formulation of the multiarmed bandit problem [82], where a set of bandits is randomly initialized at the start of each episode. Unlike conventional multiarmed bandit tasks, where reward probabilities remain fixed across episodes, this structure resets them every time. The agent must dynamically adjust its exploration and exploitation strategies to maximize long-term rewards.
I.11.10 POPGym Higher Lower
@@
I.11.11 POPGym Count Recall
@@ I.11.12 POPGym Battleship
I.11.13 POPGym Mine Sweeper
@@
I.11.14 POPGym Labyrinth Explore
@@
I.11.15 POPGym Labyrinth Escape